HelpMePlz:

What is the volume of a sphere with a diameter of 14 centimeters? Use 3.14 for pi. Round your answer to the nearest hundredth. 1077.02 cm3 1436.03 cm3 1975.43 cm3 11,494.04 cm3

Champion:

hint: volume of a sphere \(\Large V=\frac{4\pi r^3}{3}\) r=14 centimeters \(\pi\)=3.14

HelpMePlz:

thank you so much!

Champion:

you're welcome ^_^